OpenStudy (anonymous):

Using the substitution x = cos^2(x) OR OTHERWISE, find:

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 2+x }{ \sqrt{x^3(1-x)} }dx\]

OpenStudy (anonymous):

I've substituted it, but am stuck.

hartnn (hartnn):

after substitution did you get \(\huge \int -2\dfrac{2+\cos^2 x}{\cos^2x}dx\) ??

OpenStudy (anonymous):

no, but wait, lemme change a few things. I turned 1-cos^2(x) into sin^2(x), should I have done that?

hartnn (hartnn):

YES.

hartnn (hartnn):

then ?

OpenStudy (anonymous):

and the derivative of cos^2(x) can be written as -sin(2x)??

hartnn (hartnn):

no.....

OpenStudy (anonymous):

i mean, sorry, but it was cos^2(theta), not x, so we would need to change the dx to dtheta.

OpenStudy (anonymous):

no??

hartnn (hartnn):

derivative of cos^2(theta) can be written as 2 cos theta * d/dtheta (cos theta) =... ? chain rule

OpenStudy (anonymous):

which is -2sin(theta)cos(theta).

hartnn (hartnn):

yes

OpenStudy (anonymous):

but 2sinxcosx is sin2x, so why can't we make it negative?

hartnn (hartnn):

but you got how ?

OpenStudy (anonymous):

i got what how?

hartnn (hartnn):

oh...ok...its corrct, -2sin 2x but don't write it that way keep it as -2 sin x cos x

OpenStudy (anonymous):

ok, but it should be just-sin(2x), not -2sin(2x), then, i guess :P

hartnn (hartnn):

yeah :P

OpenStudy (anonymous):

ok, got that.

hartnn (hartnn):

then what gets cancelled ?

OpenStudy (anonymous):

I'm guessing the root, but idk how.

hartnn (hartnn):

\(\huge \sqrt {\cos^6x \sin^2 x}= \cos^3x \sin x \)

OpenStudy (anonymous):

OMG HOW DID I NOT SEE THAT. KILL ME. NOT LITERALLY.

hartnn (hartnn):

then one sin and one cos from both numerator and denominator will get cancelled and you will get \(\huge \int -2\dfrac{2+\cos^2 x}{\cos^2x}dx\)

OpenStudy (anonymous):

Ok, I have that. now wait, lemme see what I can do.

hartnn (hartnn):

now its a piece of cake :D :P

OpenStudy (anonymous):

split the fractions, ugh.

hartnn (hartnn):

yes.

OpenStudy (anonymous):

hold on, if x = cos^2(theta), then can we say that (x)^0.5 = cos(theta), and therefore theta = cos'(x^0.5) ??

hartnn (hartnn):

yes, but what is that needed ?

OpenStudy (anonymous):

because we have 1, and when we integrate that we get theta, and we need to give the ans in terms of x :P

hartnn (hartnn):

ohhhhh.....yes. but do you have an answer to verify ?

OpenStudy (anonymous):

nope, sorry, this kinda worksheet is like a take home test. No answers until he corrects it and I get everything wrong :P lol jk.

hartnn (hartnn):

haha..ok \(\large \cos^{-1}\sqrt x\) in the answer is pretty weird but its correct here...

hartnn (hartnn):

and so is \(\large \tan [\cos^{-1}\sqrt x]\)

OpenStudy (anonymous):

ok THANK YOU SO MUCH

hartnn (hartnn):

ok WELCOME SO MUCH ^_^

OpenStudy (anonymous):

and you better hang around, I might post a few more q's.

hartnn (hartnn):

sorry :(

OpenStudy (anonymous):

sorry??

hartnn (hartnn):

i was leaving....